### v2

# Hadamard Encoded Row Column Ultrasonic Expansive Scanning (HERCULES) with Bias-Switchable Row-Column Arrays

Abstract
Top-Orthogonal-to-Bottom-Electrode (TOBE) arrays, also known as bias-switchable row-column arrays (RCAs), allow for imaging techniques otherwise impossible for non-bias-switachable RCAs. Hadamard Encoded Row Column Ultrasonic Expansive Scanning (HERCULES) is a novel imaging technique that allows for expansive 3D scanning by transmitting plane or cylindrical wavefronts and receiving using Hadamard-Encoded-Read-Out (HERO) to perform beamforming on what is effectively a full 2D synthetic receive aperture. This allows imaging beyond the shadow of the aperture of the RCA array, potentially allows for whole organ imaging and 3D visualization of tissue morphology. It additionally enables view large volumes through limited windows. In this work we demonstrated with simulation that we are able to image at comparable resolution to existing RCA imaging methods at tens to hundreds of volumes per second. We validated these simulations by demonstrating an experimental implementation of HERCULES using a custom fabricated TOBE array, custom biasing electronics, and a research ultrasound system. Furthermore, we assess our imaging capabilities by imaging a commercial phantom, and comparing our results to those taken with traditional RCA imaging methods. Finally, we verified our ability to image real tissue by imaging a xenograft mouse model.
3D imaging, hadamard encoded readout (HERO), hadamard encoded row column ultrasonic expansive scanning (HERCULES), row-column arrays, top-orthogonal-to-bottom electrode (TOBE) arrays,
1 Introduction
Ultrasound, as a low-cost, non-invasive, accessible imaging modality can enable various diagnostic and clinical applications. Despite this flexibility, ultrasound usability is limited by the difficulty in properly assessing and estimating features while relying only on a 2-dimensional view. This creates a reliance on trained operators, as wells as operator-dependent variance in measurements.
These problems have prompted interest in 3D ultrasound imaging [1, 2, 3, 4, 5]. Volumetric imaging can also better enable other ultrasound applications such as flow imaging, elastography, and super-resolution imaging [6, 7]. A wide spectrum of information in a 3D field of view could give clinicians a comprehensive overview of organ health and operation, with a non-invasive, and accessible imaging modality.
To achieve this goal, better visualize tissue morphology and perform full organ imaging with ultrasound, there has been increased research in the development of 3D ultrasound imaging. Various methods have been used, but most methods are subject to various limitations. Mechanical scanning makes use of traditional ultrasound arrays with their proven imaging characteristics in a 2D plane, but this tends to create motion artifacts, and framerate is limited due to physical constraints of mechanical systems [8]. Fully wired matrix probes can image with high resolution and contrast, but require complex wiring and high element counts and are thus limited in size and resolution. Additionally, they often need to make use of microbeamformers which introduce imaging artifacts [1, 9].
Row-Column addressed Arrays (RCAs) are able to image in a wide field of view at high framerates, but are unable to simultaneously focus in both the azimuthal and elevational directions [10, 11, 12]. This creates asymmetrical resolution properties that can be worse than matrix probes, but they require far lower element counts. As opposed to matrix arrays whose element count scale quadratically with the size of the array, RCAs scale linearly, as they are effectively two linear arrays stacked orthogonally on top of each other. Various acquisition strategies have been used to mitigate the disadvantages of RCAs, most notably coherently compounding multiple acquisitions with different virtual sources [13, 14, 15].
Volumetric scanning with RCAs creates additional challenges. RCA imaging schemes typically try to take advantage of the benefits presented by ultrafast ultrasound techniques [16, 17, 18, 14, 19]. One of the most significant drawbacks of ultrafast ultrasound imaging with linear arrays is the lack of elevational focusing, but RCAs can mitigate this issue by transmitting with the rows of the arrays while receiving ultrasound signals with the columns, as show in Figure 1 (f) [20]. However, this creates a new problem; while one aperture can project energy in a wide field in a plane, the other will lack any focusing ability in that direction and thus limit the field of view to below the shadow of the aperture. This limits the feasibility of RCAs for clinical applications requiring limited viewing windows but large volume requirements such as cardiac and intracranial imaging [21, 22]. Diverging acoustic lenses can be used to obtain a wider field of view, but RCAs would still be limited to a cylindrical swath.
Electrostrictive transducers only act like piezoelectrics when a DC bias voltage is applied across them. This allows for parts of the array to be selectively ’turned on and off’. This property has unique applications when applied to row-column arrays, where the two linear arrays are interconnected; choosing to selectively bias the rows can enable or disable portions of each column and vice versa. Bias-sensitivity opens up the possibility of encoding additional information into transmit sequences, also known as bias-encoding. At an extreme every row-column element pair combination can potentially be addressed to create a unique aperture. Electrostrictive materials also demonstrate that the polarity of transmitted and received ultrasound signals invert depending on the polarity of the DC bias across the electrodes [23]. Combined with the general linearity of small scale ultrasound signals, this can be leveraged to achieve higher average signal-to-noise-ratios while performing bias encoding [24, 25, 26].
In a previous work we demonstrated that we can use electrostrictive bias encoding to achieve transmit encoding and create an effective synthetic transmit aperture. This effective aperture allows us to make high resolution 2D images, that have the potential to even surpass the imaging quality of matrix probes in a 2D plane [24]. We have also shown in previous work that we can perform volumetric photoacoustic imaging with RCAs by using selective biasing [27]. In this work we will build on these achievements by demonstrating that we can perform volumetric ultrasound imaging using bias-sensitive row-column arrays. Using bias encoding, we will create a virtual synthetic receive aperture reminiscent of a fully wired matrix probe. Furthermore, unlike traditional RCA imaging methods, this imaging scheme will be able to use cylindrically diverging waves to imaging beyond the shadow of the aperture of the probe. This imaging scheme is call Hadamard Encoded Row Column Ultrasonic Electronic Scanning (HERCULES).
2 Methods
2.1 Hadamard-Encoded Read-Out (HERO) Acquisition
For the follow description we will describe the rows as spanning the x-axis and varying in position along the y-axis, while the columns span the y-axis and vary in position along the x-axis as shown in Figure 1 b). Additionally, the x-axis corresponds to the lateral direction, and the y-axis corresponds to the elevational direction.
Let us model a typical row-column array by subdividing the array into a grid of ’physical elements’ delineated by the overlap of row and column electrodes. Assuming the transmit pattern is constant for every transmit event , we will call the signal at row and column .
If we receive from columns, while biasing the rows with a bias pattern selected from a Hadamard matrix , the column signals for a transmit event can be modeled as:
Parameter | Value |
Speed of Sound | 1540 m/s |
Center frequency | 6.3 MHz |
Sampling frequency | 50 MHz |
Pitch | 250 m |
Kerf | 30 m |
Number of excitation cycles | 1 |
2D array size | 128 128 |
Transmit Count | 128 |
Point Target Depth | 50 mm |
This may be written in matrix form as . After acquiring from the complete set of transmit events, the channel data from every element can then be estimated as . This procedure constitutes the Hadamard-Encoded Read-Out (HERO) acquisition scheme. This was previously used for photoacoustic readout, but not yet for volumetric ultrasound imaging [27]. Note that the row biasing patterns may alternate polarity, which for electrostrictive materials will result in an inversion of the transmit waveform. To ensure that the same transmit pattern is emitted for every transmit event, we must account for this inversion. Our approach thus inverts the polarity of the transmit waveform for rows where the biases are negative [28].
As an example, for an 128 + 128 channel TOBE array, we can use 128 transmit events to obtain the effective results of a diverging wave transmit and full synthetic receive acquisition from a fully wired 128 128 matrix array. Figure 1 illustrates an arbitrary N-transmit HERO acquisition. If we use a virtual source behind the array we can generate a diverging cylindrical wave, and as our receive aperture is a 2D array, we can focus everywhere in receive. We can repeat this after swapping the roles of the rows and columns, and image a notched pyramidal volume.
2.2 Simulation
We hypothesize that HERCULES will provide imaging characteristics that are comparable to traditional row-column array, with the potential to also result in a reduction in artifacts. Additionally, we expect that we can achieve comparable SNR when imaging a point outside the shadow of the aperture as when imaging a point under the shadow if we use a diverging wave. To this end, we will compare 3 imaging schemes along with some variations in silico. These imaging schemes are: 1) RCA imaging using a walking virtual line source (VLS) for transmission [29]. 2) RCA imaging using tilted plane wave (TPW) transmissions [30]. 3) HERCULES imaging.
We use Field II [31, 32] to simulate point spread functions for each imaging scheme. We model our RCAs as a 2D array of physical elements and apply an apodization reflecting the bias states of the intersecting row & column. Receive signals are created by adding up the signals of every element for a given row or column. For a fair comparison we utilize the same array specifications for all simulations; these parameters are listed in Table 1. Additionally, we maintain a constant number of transmits for a consistent acquisition time and energy output.
Acquisition | HERCULES Planewave (0°) | HERCULES Diverging (45°) | Virtual Line Source | Tilted Planewave |
Lateral Resolution (m) | 675 | 675 | 1713 | 625 |
Elevational Resolution (m) | 675 | 675 | 675 | 675 |
Axial Resolution (m) | 175 | 187 | 175 | 175 |
CNR | 0.6449 | 0.4094 | 0.2484 | 0.3436 |
For this comparison we evaluate various imaging parameters. We measure the lateral, elevational, & axial resolutions via the respective full-width-half-maximum (FWHM) value. To obtain an effective measure of the relative contrast to noise ratio (CNR) between the different imaging schemes, we will measure the ratio of energy within 2.5 wavelengths of the center of the PSF to the energy outside that radius. These measurements will later be further corroborated with contrast measurements on experimental data. (Figure 3) shows simulated Point-Spread-Functions (PSFs) for the aforementioned imaging schemes.
2.3 Experiment
To experimentally evaluate and compare these imaging schemes we use custom fabricated TOBE arrays (CliniSonix, Edmonton, AB, Canada) [33, 34]. These arrays can act as standard row column arrays, when one side is uniformly biased, and will thus allow for a consistent and fair comparison between standard RCA imaging methods and our bias-encoded imaging methods. Our TOBE arrays require that bias voltages be provided to individual channels independently. We will use custom biasing electronics to do this [35]. Ultrasound transmit signals will be provided by an attached Vantage 256 High Frequency Ultrasound Research System (Verasonics, Kirkland, WA, USA). This system will also be responsible for receiving ultrasound signals. DC bias signals from the biasing electronics, and AC ultrasound signals from the Vantage System will be coupled together on bias tees and sent to the array. This is shown in Figure 2. To demonstrate the structural imaging capabilities of these imaging methods we will use low-to-mid frequency arrays, suitable for large organ imaging. Specifically, we will use a low-frequency 3-6 MHz 4545 mm and a mid-frequency 4-8 MHz 3232 mm TOBE handheld probe. To calculate imaging parameters we will image an ATS 539 commercial quality assurance phantom (CIRS Inc., Norfolk, VA, USA). We seek to evaluate the resolution, CNR, and viewing angles of these different imaging methods, as well as their consistency across the imaged volume.
Compared to conventional focused imaging, ultrafast imaging suffers from a natural disadvantage; any given part of the imaged medium, is insonified to a far lesser degree, and thus the resulting image tends to suffer from lower SNR & CNR [36, 37]. To counteract this, we can lengthen the transmitted pulse, but this will potentially reduce our resolution. When imaging low contrast targets, we will use chirp pulses as these provide the long pulses we need, but can be filtered to nearly fully recover the lost resolution [38, 39, 40]. For volumetric imaging of the cyst targets we used 30s long chirps with a bandwidth matching the bandwidth of the transducer.
To further evaluate the different imaging schemes, we performed imaging ex-vivo in mice. Tumor-bearing mice were used for these experiments. Animal experiments were approved through the University of Alberta ACUC (AUP #3982 and #2994). SCID Hairless Outbred (SHO, Charles River) mice were injected with subcutaneous B16F10 mouse melanoma tumors in the hind flank. Once flank tumors had grown to a size of at least 5 mm in diameter, the hind section of the mice were then submerged in degassed water and imaged with our 8MHz TOBE arrays.
2.4 Processing
We use intensity modulated ultrasound receive data from the Vantage System, and manipulate it with the rest of our beamforming pipeline to reconstruct our images. The first step is to apply a matched filter that models our transmit waveform to allow us to recover the effective results of a single cycles transmit. We also apply the Hilbert transform to obtain the analytical signal, and improve the effects of compounding in later processing stages. For our HERCULES data we then multiply the inverse of our encoding Hadamard matrix across transmits, to recover acquisition data for every element of our 2D grid of receive elements. From this point we perform standard delay-and-sum beamforming[41] to reconstruct an intensity image. We generate low resolution images from every transmit and every receive element, and coherently compound these images together to obtain a final high resolution image. During experiments, we perform this beamforming at real-time imaging rates. We perform offline volume reconstruction with custom beamforming software and scripts to validate, and perform post-processing on our results. 2D images are normalized to their maximum values. 3D Volumes are rendered in MATLAB [42], with smoothing and thresholding.
To measure resolution we use the line targets present in the ATS 539 phantom. We cut an elevational slice from the imaged volume, and measure the lateral and axial resolution of the points in the image. Similarly, to measure contrast we use the anechoic cyst targets in the phantom. For our contrast metric we use generalized Contrast-to-Noise-Ratio (gCNR) [43] and compare histograms of the pixels inside and outside the cysts to measure their detectability. In this case we beamform rectangular volumes centred on the cyst, elongated in the y-direction, and extract concentric cylindrical regions that represent inside, and outside the cyst, without including the transition region.
3 Results
We simulated the PSF at a depth of 50 mm for a 128 128 array. In Figure 3 we see that all the imaging methods create well-resolved points in the elevational-axial plane. RCAs are susceptible to the effects of edge waves, which results in echoes below the main lobe (a-h)) [21]. The axial and elevational resolutions of the various imaging methods are nearly identical, as they are primarily dependent on the wavelength of the ultrasound waves, and the physical properties of the array. On the other hand, the there is some variation laterally; While HERCULES sees no difference compared to it’s elevational properties (i-j)), the use of tilted plane waves or virtual line sources creates significant differences (k-l)). We also measured the ratio of energy within 2.5 wavelengths of the point location, to that outside, and found that HERCULES maintained a higher inner energy ratio. This suggests that HERCULES will be less susceptible to sidelobes and be able to attain higher contrast when imaging complex structures. These results are summarized in Table 2.
In Figure 5 we validate our resolution measurements by imaging a series of line targets. Here, we obtain similar results from our simulations. We can see well resolved points in every imaging scheme in the xz plane and lines in the yz plane. In the yz plane we are unable to image beyond the shadow of the aperture, except for in the case of HERCULES 45° (b)). We additionally show how lateral resolution varies with depth in Figure 5 by transmitting on the columns. As expected, the resolution degrades as we go deeper as the relative aperture decreases. TPW is the most susceptible to this effect in its transmit plane. We characterize the imaging methods’ elevational resolution by transmitting on the rows, and we can see that there is effectively no difference between methods.
While HERCULES’ effective imaging function benefits from not needing to assume the separability of its transmit and receive apertures, its primary advantage is the ability to image cylindrical fields of view. In Figure 6 we image anechoic cysts and assess the contrast of our imaging schemes, the results of which are shown in Figure 7. We can see that all the imaging methods have comparable image quality and penetration, with a slight decrease in HERCULES 45°, but this is to be expected, as the transmitted energy is dispersed over a larger area. Secondly, the visibility of cysts, beyond the shadow of the aperture dramatically decreases, again with the exception of HERCULES 45° (Figure 6 b)). This is further corroborated quantitatively by Figure 7 d) where we obtain higher gCNR values for HERCULES 45° despite it being the lowest in a-c). The supplementary material contains flythroughs of these comparisons. It also contains a sector scan that shows how we can continuously vary the viewing region by changing the transmit delay profile. If HERCULES is used with a diverging acoustic lens, it would even be possible to image a full pyramidal volume, something that would normally only be achievable with a matrix, array; but even a matrix array cannot do this at high resolution due to channel count requirements.
Acquisition | VLS | TPW | FORCES/uFORCES | HERCULES |
References | [29] | [30] | [24, 25, 26] | N/A |
Possible with Row-Column Arrays | Yes | Yes | No | No |
Possible with Electrostrictive TOBE Arrays | Yes | Yes | Yes | Yes |
Description | – Virtual line sources emissions on columns while receiving on all rows (or vice versa). – Synthetic transmit focusing in azimuth, dynamic receive focusing in elevation (or vice verse). | – Tilted plane wave emissions with reception on rows then columns. | – Transmit Aperture encoded multi-static synthetic aperture imaging with electronically-steerable scan plane and elevation focus. | – Plane or cylindrically diverging wave emissions, reception on every element using Hadamard aperture-encoded readout (HERO). |
Implementation | – Tx on columns with walking cylindrical wave emissions, receive on rows (or vice versa). | – Tx on columns to form tilted plane wave emissions, receive on all rows, then reverse roles of rows and columns. | – Tx on rows while biasing and Rx on columns (or vice versa). Invert polarity of Rx signals when negatively biased, aperture-decode and beamform. | – Bias and Tx on columns, inverting Tx polarity when negatively biased, Rx from rows. Aperture decode to emulate fully-wired 2D array on reception. Can reverse roles of rows and columns. |
Capabilities | – 3D imaging – Seperable row- and column focusing | – 3D imaging – Seperable row- and column focusing | – Very high-quality B-scan beyond the shadow of the aperture – 2D images with transmit and receive focusing (2-way focusing) everywhere in-plane – Electronically-steerable scan plane and elevation focus | – 3D imaging beyond the shadow of the aperture – Similar image quality to VLS and TPW methods but larger FOV – Spherical receive focusing everywhere in the volume |
Limitations | – One-way elevation focusing – One-way azimuthal focusing – Imaging only beneath the shadow of the aperture | – One-way elevation focusing – One-way azimuthal focusing – Imaging only beneath the shadow of the aperture | – One-way elevation focusing – Transmit focusing limited to signal elevational focal depth | – Full pyramidal volumes not yet achieved – Unfocused transmissions lead to one-way focusing (receive focusing only) |
Hardware Needed | – Research Ultrasound Platform + RCA | – Research Ultrasound Platform + RCA | – Research Ultrasound Platform + TOBE Array + Biasing Electronics | – Research Ultrasound Platform + TOBE Array + Biasing Electronics |
We also show that HERCULES is able to effectively image real tissue. In Figure 9 we image a xenograft mouse model cadaver, with all imaging methods. A tumor can be seen on the hind leg of the mouse. In all these volumetric imaging methods sidelobes from the strong scattering against the surface of the mouse and its bones create artifacts that reduce the overall quality of the volume. These artifacts are characteristic of the volumetric imaging limitations of row-column arrays. We show a slice from the HERCULES volume (a)) as proper 3D visualization of ultrasound data is a non-trivial problem. It is difficult to make 2D arrays suitable for small animal volumetric imaging, but TOBE arrays have the potential to fill this space.
For ultrafast imaging, limiting transmit counts is important to minimize the effect of motion artifacts and allow us to perform motion compensation, as well as enable applications like flow imaging which require the collection of many volumes per second [44, 45]. All the previous images were made with 128 transmit events. With VLS & TPW we can simply reduce the amount of transmits, use sparser effective transmit apertures and create lower quality volumes. This is less trivial for HERCULES, as Hadamard matrices do not exist for an arbitrary number of elements, we will no longer be decoding the entire array. However, that is not to say that there are no other options to reduce transmit count; grouping elements into bins and using a smaller Hadamard matrix is one possible solution. Active motion compensation can also potentially be used to alleviate motion artifacts.
Figure 8 demonstrates a unique advantage of TOBE arrays; large volumes of tissues can be scanned using HERCULES, and then high quality and resolution images can be obtained using a companion imaging method Fast-Orthogonal-Row-Column-Electronic-Scanning (FORCES) [24, 25]. Table 3 summarizes the differences between TOBE imaging methods and regular RCA methods.
4 Conclusion
In this work we were able to demonstrate a new imaging method that leveraged the properties of bias-sensitive row column arrays with bias encoding to perform high resolution volumetric structural imaging at real-time rates without mechanical scanning. This method has comparable imaging quality to other row-column imaging techniques, with the additional unique benefit of being able to image beyond the shadow of the aperture. These properties help to open the way for various clinical applications such as full organ imaging, and imaging volumes through limited apertures, such as cardiovascular imaging through the ribs, and intracranial imaging. Future work could include implementation of methods to reduce the required transmits to generate a volume. Volumetric flow, super-resolution, and elastography imaging could also build on this work.